Iodine-123 is used to study thyroid gland function. This radioactive isotope breaks down in a
first-order process with a half-life of 13.1 h. What is the rate constant for the process?
1
Expert's answer
2021-10-17T11:21:31-0400

"K=\frac{ln2}{t_{\frac{1}{2}}}=\frac{0.693}{13.1}=5.29×10^{-2}hr^{-1}"